# Experimentally Testing AI-Powered Content Transformations on Student Learning

### v1

@@ 4.1. Research Hypotheses
    We hypothesize that participants will perform better on both an immediate and long-term recall assessment in the Learn Your Way condition compared to the Digital Reader condition, drawing on prior research demonstrating that transforming static content into different modalities can have a positive effect on learning outcomes (Ainsworth, 1999; Do et al., 2025).
@@
    Consistent with research that demonstrates that students experience enjoyment from AI-generated tools (Abdaljaleel et al., 2024; Do et al., 2022, 2025; Leong et al., 2024), we hypothesize that participants in the Learn Your Way condition will report more positive learning experiences compared to those in the Digital Reader condition.
4.2. Dependent Variables
All assessments and questionnaires were programmed and administered through Qualtrics444www.qualtrics.com. Full question banks can be found in the Supplementary Materials.
4.2.1. Learning Outcomes
To assess immediate recall and comprehension, participants completed a  10 minute, in-session assessment that consisted of short answer questions (SAQs), single-answer multiple choice questions (SMAQs), multi-answer multiple choice questions (MMAQs), and matching questions. The immediate recall assessment was scored out of 12 points.
@@ 4.2.1. Learning Outcomes
Both assessments were developed by pedagogical experts and targeted specific levels of Bloom’s Taxonomy in order to ensure cognitive complexity (Bloom, 1971). The assessments were therefore designed to be a good measure of content comprehension and long-term knowledge consolidation (see Supplementary Materials for all assessment questions).
In order to establish interrater reliability of SAQ scoring for both assessments, one of our pedagogy experts defined rubrics with scoring criteria and examples for each SAQ on a common 3-point scale (see Supplementary Materials for rubrics). Our pedagogy expert and a member of the research team independently scored and compared results for one third of the data to ensure alignment and adherence to the rubric guidelines. Both were blind to the experimental condition. The remainder of SAQ scoring was completed independently by the study’s pedagogy expert.
4.2.2. Learning Experience
@@
Figure 2. Overview of the experimental set-up.:
4.3. Recruitment
@@
To ensure students participating in the study had similar levels of reading comprehension, they read a short passage and completed a reading comprehension task during the recruitment process. The average score on the assessment was 6.4 out of 10, with a standard deviation of 2.3. We included students who scored 1 standard deviation above or below the mean, or those who scored a 4-9 out of 10 on the assessment. The study was conducted in adherence to Google and Bold Insight’s ethical, legal, and privacy standards for human subjects research, and our participants completed an informed consent process.
4.4. Procedure
The experiment consisted of a single 90 minute in-person lab session and a follow-up assessment shared remotely with participants three days post-session. See Figure 2 for an overview of the experimental setup.
At the beginning of the study, participants were randomly assigned to one of the two experimental conditions (Learn Your Way or Digital Reader) and were given a brief overview (three slides) that introduced the educational tool they would be using during the task (see Supplementary Materials). They were then asked to imagine they were sitting in one of their classrooms at school where their teacher asked them to learn the educational content using the provided tool and complete an assessment based on what they learned. Participants in the Learn Your Way condition selected a preloaded PDF of the learning material and indicated the grade level they wanted to learn the content in, as well as their interests (e.g., music, video games, biking, etc). All participants in the Learn Your Way condition started in the Immersive Text modality, and therefore interacted with the QuizMe feature. Participants in the Digital Reader condition opened the learning material in a PDF reader that was locally installed. To ensure uniformity across conditions, all participants were given up to 40 minutes (but required to spend at least 20 minutes) using the tool and learning the content before taking a short assessment. We report the maximum and average learning duration by educational tool condition in Table 1.
Table 1. Duration of participant tool usage for Learn Your Way and Digital Reader.
@@ 4.4. Procedure
Duration of tool use Learn Your Way Digital Reader
@@
Table 2. Participant demographics by educational tool condition, location, and grade level.
@@
After the allotted 40 minutes (or when the participant indicated they were ready to move on), participants completed the immediate assessment ( 10 minutes) and the learning experience questionnaire ( 5 minutes) via a survey. Finally, they spent 30 minutes in a qualitative, debrief interview with a skilled moderator. In these in-depth interviews, participants were asked questions about their experience using the educational tool, utility of the tool and corresponding capabilities, experience with the assessment, recommendations for improving the tool, and how the tools compare to current tools they use for learning.
Three days post-session, participants were emailed a follow-up assessment that measured long-term recall of the content they had learned during the experiment. 58 out of 60 participants completed the long-term recall assessment within the required 3-7 day window. Participants were compensated between \$115- \$160, accounting for study time, travel time, and completion of the follow-up survey.
4.5. Participants
@@
Participants were randomly assigned to one of two educational tool conditions: Learn Your Way (treatment condition) or Digital Reader (control condition). Table 2 shows a breakdown of participants’ demographics across conditions.
4.6. Data analysis and Approach
@@
To address our research questions, we analyzed the collected data at multiple levels. All data and statistical analyses were conducted in R and RStudio including the statistical package coin (Hothorn et al., 2008).
@@
First, to inform the statistical method for immediate recall and long term recall, we conducted a Shapiro-Wilk test on each dependent variable to check for the parametric assumption of a normal distribution (Shapiro and Wilk, 1965). We addressed any violations of this assumption by conducting a non-parametric alternative to the independent t-test, the Wilcoxon-Mann-Whitney test. We report the Z-statistic and subsequent effect size for each test. In addition, we conducted directional Wilcoxon-Mann-Whitney tests for our learning experience metrics.
@@
Second, we conducted several bias checks to assess whether extraneous factors may be driving the primary effects we observed. To explore whether students’ prior reading comprehension levels influenced our learning outcomes, we compared screener assessment scores for Learn Your Way participants to those for Digital Reader participants through a Wilcoxon-Mann-Whitney test on a reading comprehension task (eligible scores for participation ranging from 4-9 points). Next, to ensure the learning material (textbook chapter) struck a balance of being relatively unfamiliar to participants (as to not bias their recall scores), but still interesting enough (to encourage participant engagement in the task), we collected participants’ baseline familiarity with the learning material in two ways. First, through a screener questionnaire and again, during the in-session survey. In addition, we evaluated participant interest in the content. Content familiarity and content interest were all analyzed separately with two-sided Wilcoxon-Mann-Whitney tests, appropriately suited for ordinal data.
Finally, for our qualitative interviews, we applied Braun and Clarke’s thematic analysis method (Braun and Clarke, 2021) to generate themes that helped us understand the key patterns observed in the quantitative data in more depth. This approach enabled us to identify the commonalities and differences in students’ behavior, attitudes, and reflections through personal accounting of their learning experiences.
5. Results
5.1. Bias Checks
@@
A Shapiro-Wilk test confirmed a significant departure from normality for reading comprehension scores (N= 60, W= 0.93, p= 0.002). We therefore conducted a non-parametric Wilcoxon-Mann-Whitney test and found no significant difference in scores on the reading comprehension task between Learn Your Way (Mdn= 7; M= 6.5, SD= 1.4) and Digital Reader students (Mdn= 7; M= 6.7, SD= 1.4), Z= 0.54, p= 0.59, r= 0.07.
@@
In addition, we explored whether prior familiarity with the learning material differed by experimental condition for either the screener ratings or in-session ratings of familiarity. We found no significant difference in baseline familiarity with the learning material between Learn Your Way (Mdn= 2.5; M= 2.6, SD= 1.1) and Digital Reader conditions (Mdn= 3; M= 2.9, SD= 1.2) for screener ratings of familiarity (Z= 0.98, p= 0.34, r= 0.13). Additionally, no significant difference was detected between the Learn Your Way condition (Mdn= 3; M= 2.6, SD= 0.9) or the Digital Reader condition (Mdn= 2.5; M= 2.4, SD= 0.9) for content familiarity assessed during the in-session questionnaire (Z= 0.52, p= 0.62, r= 0.07).
In terms of interest in the topic, we also found no significant difference between Learn Your Way (Mdn= 3; M= 3.3, SD= 0.6) and Digital Reader conditions (Mdn= 3; M= 3.2, SD= 0.8), Z= 0.11, p= 0.93, r= 0.01. In fact, the majority of participants, regardless of experimental condition, rated the learning content as somewhat interesting or very interesting on a 4-point Likert scale.
@@
Lastly, we evaluated students’ perceived assessment difficulty, as it was important for the immediate recall assessment to feel like a reasonable task for high school students – in that it was not too difficult nor too easy. We found no significant difference in student ratings of assessment ease or difficulty between the Learn Your Way condition (Mdn= 4; M= 3.6, SD= 0.6) and the Digital Reader condition (Mdn= 3; M= 3.4, SD = 0.8), Z= 1.1, p= 0.27, r= 0.15.
@@
5.2. Learning Outcomes
In support of our first hypothesis (H1), we found that students in the Learn Your Way condition (M= 9.2, SD= 2.2) performed significantly better on the immediate recall assessment compared to students in the Digital Reader condition (M= 8.2, SD= 2).
In addition, analysis of long-term recall scores provides further support for H1, with Learn Your Way participants outperforming Digital Reader participants on the 3-7 day post-session assessment. Students in the Learn Your Way condition (M= 4.7, SD= 1.2) scored significantly higher on the long-term recall assessment than students in the Digital Reader condition (M= 4, SD= 1.4).
In Table 3, we further breakdown immediate and long-term recall performance, including median scores and interquartile range (IQR), and in Table 4, we show the results of statistical tests performed. Finally, we use violin and box plots to represent the overall score distributions for Learn Your Way and Digital Reader in Figure 3.
Table 3. Median (Mdn), Interquartile range (IQR), and sample size (n) of immediate recall and long-term recall assessment scores by educational tool condition. The immediate recall assessment was scored out of 12 points while the long-term recall assessment was scored out of 6 points.
@@ 5.2. Learning Outcomes
Digital Reader 8.0, 2 (30) 4.0, 2 (29)
Table 4. Summary of Wilcoxon-Mann-Whitney tests of immediate and long-term recall scores with directional hypothesis testing. Astrisks denote significant differences at p¡ .05. Effect sizes are reported as the rank-biserial correlation.
@@
Learn Your Way > Digital Reader
@@
p = 0.03*
r = 0.25
Figure 3. Violin and boxplots of immediate recall scores (left plot) and long-term recall scores (right plot) by tool condition. The immediate recall assessment was scored out of 12 points while the long-term recall assessment was scored out of 6 points.:
When looking at the differences in students’ scores in the Learn Your Way condition, students’ performance was almost identical between the immediate recall assessment (M = 9.2 / 12 points, or 77%) and long-term recall assessment (M = 4.7 / 6 points, or 78%). This result suggests that Learn Your Way students retained similar levels of information from the learning material, even after a delay of 3-7 days post-session.
@@
Table 5. Number of Learn Your Way participants who used one or more of the five learning modalities provided in the tool.
Learn Your Way
Modality Usage
@@
Table 6. Learn Your Way modality usage across participants and mean scores for the immediate recall assessment (n= 30; out of 12 points) and the long-term recall assessment (n= 29; out of 6 points). *One of these participants did not complete the long-term recall assessment, hence the drop from total n=30 during the immediate recall assessment to n=29 for the long-term recall sample.
Learn Your Way
Modality Usage
@@
Immersive Text +1 additional modality 4 10.8 (1.9) 5.3 (0.5)
Immersive Text +2 or more modalities 14* 9.3 (2.4) 4.7 (1.3)
We further split Learn Your Way participants into students that used Immersive Text only, students that used Immersive Text plus one additional learning modality and students that used Immersive Text plus two or more learning modalities (Table 6). For the following analysis, we collapsed between the latter two groups when comparing immediate assessment scores between Immersive Text + QuizMe only (n= 12; Mdn= 8.5; M= 8.6, SD= 1.8) vs. Immersive Text + any additional modality(ies) (n= 18; Mdn= 10; M= 9.6, SD= 2.3). A Wilcoxon-Mann-Whitney test revealed no significant difference between groups (Z= 1.48, p= 0.14, r= 0.27). Furthermore, while we see, directionally, that scores are higher for students that used more modalities in Learn Your Way, we did not find a significant correlation between learning modality usage and performance on the immediate recall assessment (Pearson’s r= 0.24, p= 0.21). In other words, we did not find evidence that students who used more content transformation modalities had a better performance on the immediate assessment.
Similarly, we did not find a significant difference between Immersive Text + QuizMe only (n= 12; Mdn= 4.5; M= 4.4, SD= 1.3) vs. Immersive Text + any additional modality(ies) (n= 17; Mdn= 5; M= 4.8, SD= 1.1) for long-term recall scores (Z= 0.87, p= 0.4, r= 0.16) nor did we find a significant correlation (Pearson’s r= 0.17, p= 0.38).
@@
5.3. Learning Experience
In support of our second hypothesis (H2), we observed meaningful differences when we analyzed students’ reported experience with the educational tools. Students in the Learn Your Way condition (Mdn= 4; M= 3.6, SD= 0.6) found the tool to be more useful for learning the educational content compared to students in the Digital Reader condition (Mdn= 3; M= 3.1, SD= 0.7). Specifically, a Wilcoxon-Mann-Whitney test revealed the Learn Your Way tool was rated significantly higher than Digital Reader (Z= 3.08, p= 0.001, r= 0.4).
Furthermore, we can see how the Learn Your Way tool positively impacted students’ learning experience as compared to students using the Digital Reader tool (as shown in Table 7). Specifically, students in the Learn Your Way condition reported higher levels of agreement compared to students in the Digital Reader condition across all of our experiential measures.
Table 7. Mean agreement ratings for learning experience statements shown by educational tool condition in middle two columns. Summary of directional Wilcoxon-Mann-Whitney tests shown in final column. Bolded values denote significant differences at p¡ .05. Effect sizes are reported as the rank-biserial correlation.

### v2

@@ 5.1 Research Hypotheses
    We hypothesize that participants will perform better on both an immediate and long-term recall assessment in the Learn Your Way condition compared to the Textbook condition, drawing on prior research demonstrating that transforming static content into different modalities and self-directed learning can have a positive effect on learning outcomes (AINSWORTH1999; do2025; Deci2000).
@@
    Consistent with research that demonstrates that students experience enjoyment from AI-generated tools (Abdaljaleel2024; Do2022; do2025; Leong2024), we hypothesize that participants in the Learn Your Way condition will report more positive learning experiences compared to those in the Textbook condition.
5.2 Dependent Variables
All assessments and questionnaires were programmed and administered through Qualtrics333www.qualtrics.com. Full question banks can be found in the Supplementary Materials.
5.2.1 Learning Outcomes
To assess immediate recall and comprehension, participants completed a $\sim$10 minute, in-session assessment that consisted of short answer questions (SAQs), single-answer multiple choice questions (SMCQs), multi-answer multiple choice questions (MMCQs), and matching questions. The immediate recall assessment was scored out of 12 points.
@@ 5.2.1 Learning Outcomes
Both assessments were developed by pedagogical experts and targeted specific levels of Bloom’s Taxonomy in order to ensure cognitive complexity (bloom1971taxonomy). The assessments were therefore designed to be a good measure of content comprehension and long-term knowledge consolidation (see Supplementary Materials for all assessment questions).
In order to establish consistency of SAQ scoring across both assessments, one of our pedagogy experts pre-defined rubrics with scoring criteria and examples for each SAQ on a 3-point scale. All SAQ responses were evaluated solely on students’ demonstrated knowledge of the source material (ranked from 0-Fragmentary to 3-Demonstrating), excluding any consideration of students’ grammar skills or writing proficiency in this evaluation (see Supplementary Materials for complete rubrics). Our pedagogy expert and a member of the research team independently scored and compared results for one third of the data to ensure alignment and adherence to the rubric guidelines. Both were blind to the experimental condition. The remainder of SAQ scoring was completed independently by the study’s pedagogy expert.
5.2.2 Learning Experience
@@
Figure 2: Overview of the experimental set-up.
5.3 Recruitment
@@
To ensure students participating in the study had similar levels of reading comprehension, they read a short passage and completed a reading comprehension task during the recruitment process. The average score on the assessment was 6.4 out of 10, with a standard deviation of 2.3. We included students who scored one standard deviation above or below the mean, or those who scored a 4-9 out of 10 on the reading assessment.
5.4 Procedure
The experiment consisted of a single 90 minute in-person lab session and a follow-up assessment shared remotely with participants three days post-session. See Figure 2 for an overview of the experimental setup. The study was conducted in adherence to Google and Bold Insight’s ethical, legal, and privacy standards for human subjects research, and our participants completed an informed consent process.
At the beginning of the study, participants were randomly assigned to one of the two experimental conditions (Learn Your Way or Textbook) and were given a brief overview (three slides) that introduced the educational tool they would be using during the task (see Supplementary Materials). They were then asked to imagine they were sitting in one of their classrooms at school where their teacher asked them to learn the educational content using the provided tool and complete an assessment based on what they learned. Participants in the Learn Your Way condition selected a preloaded PDF of the learning material and indicated the grade level they wanted to learn the content in, as well as their interests (e.g., music, video games, biking, etc). All participants in the Learn Your Way condition started in the Immersive Text modality, and therefore interacted with the QuizMe feature. Participants in the Textbook condition viewed the chapter in a web-based PDF reader. To ensure uniformity across conditions, all participants were given up to 40 minutes (but required to spend at least 20 minutes) using the tool and learning the content before taking a short assessment. We report the maximum and average learning duration by educational tool condition in Table 1.
Table 1: Duration of participant tool usage for Learn Your Way and Textbook.
@@ 5.4 Procedure
Duration of tool use Learn Your Way Textbook
@@
After the allotted 40 minutes (or when the participant indicated they were ready to move on), participants completed the immediate assessment ($\sim$10 minutes) and the learning experience questionnaire ($\sim$5 minutes) via a digital survey. Finally, they spent 30 minutes in a qualitative, debrief interview with a skilled moderator. In these in-depth interviews, participants were asked questions about their experience using the educational tool, utility of the tool and corresponding capabilities, experience with the assessment, recommendations for improving the tool, and how the tools compare to current tools they use for learning.
Three days post-session, participants were emailed a follow-up assessment that measured long-term recall of the content they had learned during the experiment. 58 out of 60 participants completed the long-term recall assessment within the required 3–7 day window. Participants were compensated between \$115–\$160, accounting for study time, travel time, and completion of the follow-up survey.
Table 2: Participant demographics by educational tool condition, location, and grade level.
@@
5.5 Participants
@@
Participants were randomly assigned to one of two educational tool conditions: Learn Your Way (treatment condition) or Textbook (control condition). Table 2 shows a breakdown of participants’ demographics across conditions.
5.6 Data Analysis and Approach
@@
To address our research questions, we analyzed the collected data at multiple levels. All data and statistical analyses were conducted in R and RStudio including the statistical package coin for our non-parametric tests (Hothorn2008) and package mgcv for the generalized additive models (wood2017gam).
@@
First, to inform the statistical method for immediate recall and long-term recall, we conducted a Shapiro-Wilk test on each dependent variable to check for the parametric assumption of a normal distribution (Shapiro1965). We addressed any violations of this assumption by conducting a non-parametric alternative to the independent t-test, the Wilcoxon-Mann-Whitney test. Specifically, we conducted one-sided Wilcoxon-Mann-Whitney tests on both immediate recall assessment scores and long-term recall assessment scores, in line with our first research hypothesis (H1). To address our second hypothesis (H2), stating that Learn Your Way participants will report more positive learning experiences than Textbook participants, we conducted directional Wilcoxon-Mann-Whitney tests for each of our learning experience metrics. Additionally, to control for the inflation of Type I error rate due to multiple comparisons, $p$-values were adjusted for the learning experience metrics using the Benjamini-Hochberg procedure to maintain a false discovery rate (FDR) of .05 (benjamini1995controlling). For all statistical tests, we report the Z-statistic, p-value, adjusted p-value (where applicable), and subsequent effect size r.
Second, as an exploratory analysis, we investigated immediate recall and long-term recall performance for Learn Your Way participants as a function of learning modality usage. We report mean values and directional usage patterns for three levels of modality usage: Immersive Text only, Immersive Text plus one additional modality, and Immersive Text plus 2 or more learning modalities utilized by Learn Your Way participants. To increase sample and our subsequent power to detect differences, we combined the latter two usage types for our statistical analyses including two-sided Wilcoxon-Mann-Whitney tests and Pearson’s correlation tests on both immediate recall and long-term recall scores.
@@
In addition, we conducted several bias checks to assess whether extraneous factors may be driving the primary effects we observed. To explore whether students’ prior reading comprehension levels influenced our learning outcomes, we compared screener assessment scores for Learn Your Way participants to those for Textbook participants through a two-sided Wilcoxon-Mann-Whitney test on a reading comprehension task (eligible scores for participation ranging from 4–9 points). Next, to ensure the learning material struck a balance of being relatively unfamiliar to participants (as to not bias their recall scores), but still interesting enough to encourage participant engagement in the task, we had them rate their level of interest and familiarity with the content. Participants’ baseline familiarity was evaluated in two ways: first, through a screener questionnaire and again, during the in-session survey. Both content familiarity and interest ratings were analyzed using two-sided Wilcoxon-Mann-Whitney tests, appropriately suited for ordinal data. For our final bias check, we conducted a two-sided Wilcoxon-Mann-Whitney test on students’ perceived assessment difficulty, rated on a standard five-point scale from “very difficult” to “very easy.”
To further ensure that the observed learning outcomes were not driven by procedural confounds, we modeled the influence of learning duration (transformed into seconds) and delay duration (within the 3–7 day window) on assessment scores. Specifically, we employed semi-parametric generalized additive models (GAMs) to 1) allow for potential non-linear dependencies between learning duration and recall performance, and 2) realistically model the non-linear memory decay typically associated with the forgetting curve (ebbinghaus1913memory). Within these models, smoothing terms were applied to both continuous predictors – learning duration and delay duration – nested by experimental condition, to allow the models to optimally fit both linear and non-linear trends. The complexity of these smooth terms was evaluated using effective degrees of freedom (edfedf), where $edf=1$ corresponds to a linear relationship and $edf>1$ indicates increasing non-linearity (wood2017gam).
Finally, for our qualitative interviews, we applied Braun and Clarke’s thematic analysis method (braun2021thematic) to generate themes that helped us understand the key patterns observed in the quantitative data in more depth. We chose this approach because it is widely used in other HCI interview-based studies (Johnson2022; Wong2025). A data collection tool was created in Google Sheets to capture participant behavior and interview data. During each session, the moderator used a paper guide to facilitate the interview and take notes while a notetaker entered participant responses in the data collection tool. Post-session, the moderator and notetaker reviewed the study data for accuracy and completeness. As needed, session videos and transcripts were also reviewed to confirm study data accuracy. The study team debriefed daily to review, discuss, and refine ongoing themes from the interviews. Each week, the team also met to discuss the study data and further align on trends and themes observed. The team met to refine the analysis until data saturation was achieved (hennink2017code) resulting in a final set of themes focused on differences in students’ behavior, attitudes, and reflections on their learning experiences.
6 Results
6.1 Bias Checks
@@
A Shapiro-Wilk test confirmed a significant departure from normality for reading comprehension scores ($N=60,W=0.93,p=.002$). We therefore conducted a non-parametric Wilcoxon-Mann-Whitney test and found no significant difference in scores between Learn Your Way ($Mdn=7;M=6.5,SD=1.4$) and Textbook students ($Mdn=7;M=6.7,SD=1.4$) on the reading comprehension task ($Z=0.54$, $p=.588$, $r=.07$).
@@
In addition, we explored whether prior familiarity with the learning material differed by experimental condition for either the screener ratings or in-session ratings of familiarity. We found no significant difference in baseline familiarity with the learning material between Learn Your Way ($Mdn=2.5$; $M=2.6$, $SD=1.1$) and Textbook conditions ($Mdn=3$; $M=2.9$, $SD=1.2$) for screener ratings of familiarity ($Z=0.98$, $p=.337$, $r=.13$). Additionally, no significant difference was detected between the Learn Your Way condition ($Mdn=3$; $M=2.6$, $SD=0.9$) or the Textbook condition ($Mdn=2.5$; $M=2.4$, $SD=0.9$) for content familiarity assessed during the in-session questionnaire ($Z=0.52$, $p=.616$, $r=.07$). We also found no significant difference between the Learn Your Way ($Mdn=3$; $M=3.3$, $SD=0.6$) and Textbook conditions ($Mdn=3$; $M=3.2$, $SD=0.8$) in terms of students’ reported interest in the chapter topic ($Z=0.11$, $p=.934$, $r=.01$). In fact, the majority of participants, regardless of experimental condition, rated the learning content as “somewhat interesting” or “very interesting” on a 4-point Likert scale.
@@
Lastly, we evaluated students’ perceived assessment difficulty, as it was important for the immediate recall assessment to feel like a reasonable task for high school students – in that it was not too difficult nor too easy. We found no significant difference in student ratings of assessment ease or difficulty between the Learn Your Way condition ($Mdn=4$; $M=3.6$, $SD=0.6$) and the Textbook condition ($Mdn=3$; $M=3.4$, $SD=0.8$) after conducting a two-sided Wilcoxon-Mann-Whitney test ($Z=1.15$, $p=.268$, $r=.15$).
@@
6.2 Learning Outcomes
In support of our first hypothesis (H1), we found that students in the Learn Your Way condition ($Mdn=9.5$; $M=9.2$, $SD=2.2$) performed significantly better on the immediate recall assessment compared to students in the Textbook condition ($Mdn=8$; $M=8.2$, $SD=2$).
In addition, analysis of long-term recall scores provides further support for H1, with Learn Your Way participants outperforming Textbook participants on the 3-7 day post-session assessment. Students in the Learn Your Way condition ($Mdn=5$; $M=4.7$, $SD=1.2$) scored significantly higher on the long-term recall assessment than students in the Textbook condition ($Mdn=4$; $M=4$, $SD=1.4$).
Table 3: Median (Mdn), Interquartile range (IQR), and sample size (n) of immediate recall and long-term recall assessment scores by educational tool condition. The immediate recall assessment was scored out of 12 points while the long-term recall assessment was scored out of 6 points.
@@ 6.2 Learning Outcomes
Textbook 8.0, 2 (30) 4.0, 2 (29)
In Table 3, we further breakdown immediate and long-term recall performance, including median scores and interquartile range (IQR), and in Table 4, we show the results of statistical tests performed. Finally, we use violin and box plots to represent the overall score distributions for Learn Your Way and Textbook in Figure 3.
Table 4: Summary of Wilcoxon-Mann-Whitney tests of immediate and long-term recall scores with directional hypothesis testing. Asterisks denote significant differences at p<<.05*. Effect sizes are reported as the rank-biserial correlation.
@@
Learn Your Way >
Textbook
@@
$p$ = .03*
$r$ = .25
Figure 3: Violin and boxplots of immediate recall scores (left plot) and long-term recall scores (right plot) by tool condition. The immediate recall assessment was scored out of 12 points while the long-term recall assessment was scored out of 6 points.
To ensure that performance outcomes were not driven by the duration of time students spent learning the content, we modeled the relationship between educational tool condition and learning duration on recall scores for both immediate and long-term recall assessments. Specifically, we employed a generalized additive model (GAM) to account for potential non-linear relationships between learning duration and assessment performance.
For immediate recall scores, the model explained 13.4% of the total deviance. The non-parametric smooth terms for learning duration were not significant for either Learn Your Way ($edf=1$, $F=0.48$, $p=.49$) or Textbook ($edf=1.82$, $F=1.43$, $p=.283$). Similar results were observed for long-term recall scores, where the model explained 9.1% of the total deviance. Again, the smooth terms for learning duration were not significant for Learn Your Way ($edf=1$, $F=1.53$, $p=.222$) or the Textbook condition ($edf=1$, $F=0.15$, $p=.703$). Collectively, these findings suggest that learning duration did not meaningfully influence performance outcomes on either the immediate or long-term recall assessment, regardless of the experimental condition.
Finally, to account for variations in the retention interval across students, we employed another GAM to model the relationship between length of delay (3–7 days) and long-term recall performance. The model explained 8.2% of the total deviance. The non-parametric smooth terms for delay duration were not significant for Learn Your Way ($edf=1$, $F=1.09$, $p=.3$) or Textbook ($edf=1$, $F=0.02$, $p=.903$), demonstrating that the specific length of delay did not significantly influence long-term recall within our required follow-up window of 3–7 days.
Regardless of experimental condition, assessment performance remained remarkably stable across time. Specifically, the Learn Your Way participants achieved nearly identical percentage scores on the immediate recall assessment ($M=9.2/12$; 77%) and the long-term recall assessment ($M=4.7/6$; 78%). Similarly, Textbook participants demonstrated a slight decrease of only one percentage point from immediate recall ($M=8.2/12$; 68%) to long-term recall ($M=4/6$; 67%). These results suggest that students retained the vast majority of acquired information, with negligible decay occurring across the 3–7 day retention interval. Consequently, the performance advantage observed for Learn Your Way during the immediate assessment was maintained through the long-term recall phase.
@@
Table 5: Number of Learn Your Way participants who used one or more of the five core learning modalities provided in the tool.
Learn Your Way Modality Usage
@@
Table 6: Learn Your Way modality usage across participants and mean scores for the immediate recall assessment ($n=30$; out of 12 points) and the long-term recall assessment ($n=29$; out of 6 points).
*One of these participants did not complete the long-term recall assessment, hence the drop from total n=30 during the immediate recall assessment to n=29 for the long-term recall sample.
Learn Your Way Modality Usage
@@
Immersive Text +1 modality 4 10.8 (1.9) 5.3 (0.5)
Immersive Text +2-4 modalities 14* 9.3 (2.4) 4.7 (1.3)
We further split Learn Your Way participants into students that used Immersive Text only, students that used Immersive Text plus one additional learning modality and students that used Immersive Text plus two or more learning modalities (Table 6). For the following analysis, we collapsed between the latter two groups when comparing immediate assessment scores between Immersive Text + QuizMe only ($n=12$; $Mdn=8.5$; $M=8.6$, $SD=1.8$) vs. Immersive Text + any additional modality(ies) ($n=18$; $Mdn=10$; $M=9.6$, $SD=2.3$). A two-sided Wilcoxon-Mann-Whitney test revealed no significant difference between groups ($Z=1.48$, $p=.144$, $r=.27$). Furthermore, while we see, directionally, that scores are higher for students that used more modalities in Learn Your Way, we did not find a significant correlation between learning modality usage and performance on the immediate recall assessment (Pearson’s $r=.24$, $p=.206$). In other words, we did not find evidence that students who used more content transformation modalities had a better performance on the immediate assessment. This suggests that the pedagogical value of the system may reside in the quality of engagement with a chosen modality rather than the breadth of modalities accessed.
Similarly, we did not find a significant difference between Immersive Text + QuizMe only ($n=12$; $Mdn=4.5$; $M=4.4$, $SD=1.3$) vs. Immersive Text + any additional modality(ies) ($n=17$; $Mdn=5$; $M=4.8$, $SD=1.1$) for long-term recall scores ($Z=0.87$, $p=.396$, $r=.16$) nor did we find a significant correlation (Pearson’s $r=.17$, $p=.380$).
@@
6.3 Learning Experience
Table 7: Mean agreement ratings for learning experience statements shown by educational tool condition in middle two columns. Summary of directional Wilcoxon-Mann-Whitney tests shown in final column, including adjusted p-values after multiple comparison corrections are applied. Asterisks denote significant differences at p<< .05* (or .01** and .001***). Effect sizes are reported as the rank-biserial correlation.